OpenStudy (anonymous):

multiply. 3sqrty^4*3sqrt81t^5

jhonyy9 (jhonyy9):

9sqrt(81y4t5)=9sqrt(9*9y4t4t)=9*9*y2*t2sqrt(t)=81y2t2sqrt(t)